I got this wrong, can you please help me. Solve the following system by addition.If a unique solution does not exist,state whether the system is inconsistent or dependent.
Section 8.2 problem 20:
x+ 5y =10
-2x -10y = -20

Answer by stanbon(75887)

1) x+ 5y =10
2) -2x -10y = -20
Multiply #1 equation by 2 to get
1) 2x+10y=20
Add this to equation #2 to get:
0+0=0
This means the two equation are really the same.
Cheers,
Stan H.